A disk with diameter of 60 cm speeds up from 20 rad/s to 40 rad/s in 5 seconds. (a) How many revolutions will the disk go through during that time period? (b) What is the average angular acceleration of the disk? (c) What is the average linear acceleration of the disk?


1
Expert's answer
2021-11-23T12:31:07-0500

(a)-(b) Let's first find the average angular acceleration of the disk:


α=ωfωit=40 rads20 rads5 s=4 rads2.\alpha=\dfrac{\omega_f-\omega_i}{t}=\dfrac{40\ \dfrac{rad}{s}-20\ \dfrac{rad}{s}}{5\ s}=4\ \dfrac{rad}{s^2}.

Then, we can find the angular displacement of the disk:


θ=ωit+12αt2,\theta=\omega_it+\dfrac{1}{2}\alpha t^2,θ=20 rads×5 s+12×4 rads2×(5 s)2=150 rad.\theta=20\ \dfrac{rad}{s}\times5\ s+\dfrac{1}{2}\times4\ \dfrac{rad}{s^2}\times(5\ s)^2=150\ rad.

Finally, we can find the number of revolutions that the disk will go through during 5 seconds:


n=θ2π radrev=150 rad2π radrev=24 rev.n=\dfrac{\theta}{2\pi\ \dfrac{rad}{rev}}=\dfrac{150\ rad}{2\pi\ \dfrac{rad}{rev}}=24\ rev.

(c) We can find the average linear acceleration of the disk as follows:


a=rα=0.3 m×4 rads2=1.2 ms2.a=r\alpha=0.3\ m\times4\ \dfrac{rad}{s^2}=1.2\ \dfrac{m}{s^2}.